### v1

# Experimental neuromorphic computing based on quantum memristor

Abstract
Machine learning has recently developed novel approaches, mimicking the synapses of the human brain to achieve similarly efficient learning strategies. Such an approach retains the universality of standard methods, while attempting to circumvent the excessive requirements, which hinder their scalability. In this context, quantum (or quantum inspired) algorithms may bring further enhancement. However, high-performing neural networks invariably display non-linear behaviours, which is diametrically opposed to the linear evolution of closed quantum systems. We propose a solution to this issue and report the first neuromorphic architecture based on a photonic quantum memristor. In detail, we show how the memristive feedback loop enhances the non-linearity and hence the performance of the algorithm. We benchmark our model on two tasks, i.e. non-linear function and time series prediction. In these cases, we highlight the essential role of the quantum memristive element and demonstrate the possibility of using it as a building block in more sophisticated networks.
I Introduction
In the last decades, machine learning has fundamentally changed the approach to solving a wide variety of problems, from everyday life to scientific scopes [1, 2, 3, 4, 5, 3, 6, 7]. In this context, artificial neural networks (ANN) have proven to be a pivotal tool, especially for tackling tasks where information needs to be extracted from high-dimensional data [8]. Nevertheless, a growing number of parameters must be trained, with correspondingly larger training data, as the tasks get more complex, resulting in the so-called memory bottleneck [9]. This sparked the development of more dedicated hardware, i.e. neuromorphic architectures, which mimic the human brain [10]. The main advantage of these models is their minimal learning requirements, which circumvents the aforementioned difficulty. Acting as the poster-child of neuromorphic components is the memrsitor, a passive circuital component that was postulated in 1971 [11] and demonstrated in 2008 [12]. Its main characteristic is that it retains memory of its past states, in the form of a hysteresis that is similar to neural synapses [13, 14, 10, 15, 16].
Figure 1: Neuromorphic computing based on quantum memristor. (a) Non-linear activation functions in neural networks are crucial for learning input/output correlations. Indeed, linear functions limit the expressiveness of the model, not allowing for nuanced decisions. (b) The basic principle of a quantum memristor is a tunable Mach-Zehnder interferometer, where its internal phase ($\theta$) is varied, according to the output registered at one of its output modes, through a feedback loop. The investigated hybrid quantum/classical models are implemented through three parts: input encoding (quantum), non-linear reservoir (quantum) and one-layered ANN (classical). In our work, the encoding of classical variables in a one qubit state exploits the amplitude of the component parallel to $|0\rangle$, i.e. $x\rightarrow\sqrt{x}|0\rangle+\sqrt{1-x}|1\rangle$ and $x|0\rangle+\sqrt{1-x^{2}}|1\rangle$. The reservoir consists in a unitary operation, composed by the product of different rotations. These rotations allow us to achieve a non-linear transformation of the input, due to the encoding that is non-linear in the density matrix of the quantum state. The loop enhances the achievable non-linearity, while implementing a short-term memory. The non-linear reservoir is then followed by a classical ANN, which is the only part of the model that is trained, producing a weighted sum of the reservoir outputs. The two addressed tasks are the prediction of a smooth non-linear function and of a random time series.
Figure 2: Experimental setup. (a) The experimental apparatus consists of a source of single photon pairs, which is pumped with a laser at 775 nm and generates degenerate photon pairs at 1550 nm. One photon is routed to the 3 input/output mode photonic integrated circuit (PIC) and the other photon is ancillary. Let us note that this circuit is implemented following the universal Reck architecture [17]. At the outputs of the PIC, photon counts are registered by superconducting nanowire single photon detectors. After the full calibration of the chip, phases are set and modulated by applying a suitable voltage to the resistive elements, that act as thermal phase shifters. The unitary operation is updated, conditioned on the statistics of previous runs. The output distribution is then fed into a classical ANN, which is the only part of the machine learning model that is trained. (b) The PIC consists in 3 Mach-Zehnder interferometers and 5 tunable phases. The $R_{i}$ indicate the resistive elements that serve as tunable thermal phase shifters, where $R_{3}$ is updated conditioned on the output of the third mode, implementing a memristive behaviour. Let us note that the state obtained tracing out the third output still preserves coherence (see section II of the Supplemental Material). The input is encoded in one qubit, through a unitary transformation $\textbf{U}_{enc}$. The unitary transformations $\textbf{U}_{1}$ and $\textbf{U}_{2}$ are hyperparameters of our model. Then, $\textbf{U}_{mem}$ amounts to the action of the memristor and is varied depending on previous outcomes. Also the feedback rule, according to which this happens, is a hyperparameter.
A related but somewhat orthogonal paradigm shift in computer science was introduced by quantum computing. The interest in this field stems, first, from the possibility of tackling non-classical processing tasks directly related to the investigation of quantum effects and, second, from the promise of outperforming standard algorithms for particular problems [18, 19, 20, 21]. However, for the latter, the only proven advantages were shown either for tasks which are out of reach of even near term state-of-the-art quantum computers or which have no known applications [22]. Hence, a research line that has attracted a lot of interest arises from the combination of machine learning and quantum computation, i.e. Quantum Machine Learning (QML) [23, 24]. This is fueled by the hope that a computational advantage can be found on practically relevant tasks even when using relatively small scale quantum information processing. Furthermore, photonic apparatuses exhibit vastly lower energy consumption, compared to standard electrical ones [25, 26], indicating the possibility of far more efficient computational platforms. However, in this context, the main difficulty lies in the deterministic implementation of non-linear operations (crucial for the learning process, see Fig. 1a), since the evolution of closed quantum systems is intrinsically unitary (and therefore linear). Moreover, the introduction of interactions with the environment results in the reduction of the coherence of the overall system which, in turn, leads to a processing power that approaches the one of classical information processing. This is the reason for which employing Fock states for QML algorithms is challenging, although they can grant enhanced expressivity and they allow reaching higher-dimensional Hilbert spaces [27, 28, 29] with respect to (weakly)-coherent light inputs [30, 31, 32, 33].
In this work, we propose a solution to this apparent conundrum and we implement the first instances of a neuromorphic computing protocol, i.e. reservoir computing (RC), on a photonic platform. We achieve this by exploiting single photon states manipulated through an integrated circuit [34]. The model may be seen as a building block for more complex networks and is composed of a non-linear fixed reservoir (equipped with classical memory), followed by a one-layered linear ANN, which is the only part that requires a training phase (see Fig. 1b). Our aim is to implement a quantum/classical hybrid algorithm, where the non-linear element is generated by a quantum system, while the ANN is classical. To achieve this goal we leverage a particular choice for the input encoding and the exploitation of a novel device: the photonic quantum memristor [35]. This element has the two-fold effect of enhancing the non-linearity of the model, while implementing a classical memory. Such a device is based on a Mach-Zehnder interferometer equipped with a feedback loop, and its behaviour is proven to be memristive. However, its potential use for machine learning tasks was only theorized and its role was not deeply investigated, as the studied task, the recognition of MNIST hand-written digits [36], can be tackled effectively through linear regression and other simple estimators. Here, we go beyond the previous study and apply this hardware to two machine-learning tasks where non-linearity is a crucial factor and cannot be omitted without compromising performance, highlighting the pivotal role of the quantum reservoir. In both of the considered tasks, we show that the dynamics of the photonic quantum memristor enhances the performance, with respect to the case where no feedback loop is implemented. Hence, the present work shows how a (photonic) quantum system can tackle tasks that were addressed with similar classical architectures [14], but with substantially less resources, paving the way to resource efficient machine learning models and to the possibility of exploring neuromorphic architectures applied to quantum tasks.
II Quantum memristor-based neuromorphic models
The photonic quantum memristor can be modeled as a tunable Mach-Zehnder interferometer, as depicted in Fig. 2. Then, its internal phase is updated according to a feedback rule that depends on the measurement outcome at one output mode, affecting the temporal dynamics of the device. One of its main features consists in the fact that, when tracing out the mode used for the feedback loop, the remaining reduced state may still be partially coherent relative to the joint state, as detailed in section II of the Supplementary Material. Consequently, further quantum information processing is possible on the output.
We want to exploit this feedback mechanism to implement non-linear operations on the input state and utilize a short-term memory, to achieve a proof-of-principle quantum reservoir computing (QRC). In detail, we encode our classical inputs in quantum states that are then injected into a quantum reservoir. This usually consists of randomly connected nodes, whose internal time-based state make up its intrinsic memory and the historical non-linearity that is necessary for the learning process (see the section I of the Supplemental Material for further details). In our implementation, one physical node realized through a Mach-Zehnder interferometer (MZI), which in concert with a feedback loop acts like a quantum memristor. Then, the randomness is given by unitary operations before and after the memristor itself. Hence, the quantum part of our model consists in the encoding of the classical data and reservoir. In particular, we exploit one-qubit states, encoded in the path degree of freedom of a single photon state, with a dual-rail scheme (see Fig.  2b). The reservoir is then composed of a sequence of three unitaries over three spatial modes, i.e. $\textbf{U}_{1}$, $\textbf{U}_{mem}$ and $\textbf{U}_{2}$. At the end, to build a proper reservoir computing scheme, a single trainable feedforward neural network follows.
Figure 3: Non-linear function prediction results. The blue dots show the experimental results for the prediction of the power function $f_{t}(x)=x^{n}$ for $n=4$ (indicated by the solid black curve). The orange ones show the results for the same task, but without feedback loop. Inputs lower than 0.9 are used for training the model, while the other ones are used for test. The case with no feedback loop is on average 34% less accurate than the one exploiting the memristor. Let us note that, to have a fair comparison, hyperparameter were optimized separately for the two cases. The dashed black line represents the case without the physical reservoir, i.e. where the inputs $x$ are directly injected to the ANN. The loss function is defined as the mean squared error MSE between the target function $f_{t}(x)=y_{t}$ and the test data $y^{k}_{test}$, i.e. $y_{t}-y_{test}$. The inset shows the MSE for power function with $n=3,4,5,6$ in the QRC case (blue) and with no feedback loop case (orange).
Figure 4: NARMA task results. (a) Sequence of inputs of the NARMA function, which are generated by sampling each $x$ independently from a uniform distribution between 0 and $\frac{1}{2}$. (b)-(d) Sequence of outputs $y$ (black), evaluated according to Eq. (3), together with the experimental prediction of our memristor-based model (blue); with no feedback loop (orange) and no reservoir (green). (e) Absolute difference between the true and predicted output. The adopted dataset contains 1000 points. Out of those, the first 20 correspond to the so-called washout, i.e. the time required to the model to be properly initialized, the training process is done on the following 480 points and then the validation is performed on the rest of the dataset. However, for the sake of clarity, the plot shows only 200 points of the test. The corresponding MSE are $5.73\times 10^{-5}$, $1.90\times 10^{-4}$ and $2.47\times 10^{-4}$. (f) Histogram displaying the absolute differences between true and predicted outputs (respectively in blue, orange and green).
We test our model for two tasks, that rely on the two key features of our photonic platform, i.e. non-linearity and short-term memory. These are a non-linear function prediction/reproduction, where we exploit mostly the non-linearity of the model, and a time series prediction (Non-linear Auto Regressive Moving Average [37, 38]), where short-term memory plays a pivotal role.
For the first task, the functions we aim at reproducing and then extrapolating are monomials:
$f(x)=x^{n},\ n\in\mathbb{Z}$ (1)
this choice is motivated by the interest in practically testing the maximal achievable non-linear behaviour of our model and the role of the feedback loop towards its enhancement.
We encode the inputs as follows: $x\rightarrow\sqrt{x}|0\rangle+\sqrt{1-x}|1\rangle$. Note that this encoding is linear when performing a projective measurement in the eigenbasis of the Pauli operator $\sigma_{Z}$ [39], as $Tr(\sigma_{Z}\rho_{x})=\sum_{i=0,1}(-1)^{i}p_{i}=2x-1$, where $p_{i}$ is the probability of getting the outcome $i$. In contrast, measurements in the $\sigma_{X}$ basis give an expectation value that is non-linear in $x$, $Tr(\sigma_{X}\rho_{x})=\sqrt{x}\sqrt{1-x}$. This non-linearity can be enhanced by a proper choice of the rotations $\textbf{U}_{1}$ and $\textbf{U}_{2}$. Then, we pick a linear feedback rule which modifies $\textbf{U}_{mem}$. In particular, the internal variable of the memristor $R_{t}$, which can be seen as the reflectivity of a beam-splitter, is updated via:
$R_{t}=\frac{1}{m}\sum_{T=t-m}^{t}R_{T}=\frac{1}{m}\sum_{T=t-m}^{t}(ap_{2T}+b)$ (2)
where $m$ is the memory extent, $p_{2T}$ is the probability of detecting a photon in the update mode and the internal phase of the MZI implementing $\textbf{U}_{mem}$ is given by $2\times arccos(\sqrt{R})$. The rotations $\textbf{U}_{1}$ and $\textbf{U}_{2}$, along with $a$ and $b$, are optimizable hyperparameters (see section IA of the Supplementary Material for further details).
The goal of the second task is to predict the output of a non-linear dynamical system, introduced in [38, 37], called NARMA. This was already widely used as a benchmark for classical [40, 41] and quantum [42, 43] RC models. The recurrence relation of the system is described by the following equation:
$y_{k+1}=0.4y_{k}+0.4y_{k}y_{k-1}+0.6x_{k}^{3}+0.1,$ (3)
where $x_{k}$ and $y_{k}$ is the input and the output, respectively, at time $k$.
From Eq. (3), it is visible that the output $y_{k+1}$ depends on past two outputs $y_{k}$ and $y_{k-1}$.
The time series $\bm{x}=\{x_{i}\}_{i=1}^{N}$ is generated by sampling each $x_{i}$ independently and uniformly in the interval $U(0,\frac{1}{2})$, and then outputs $\bm{y}=\{y_{i}\}_{i=1}^{N}$ come sequentially from Eq. (3). For this task, we are using the amplitude encoding $x_{i}\mapsto x_{i}\ket{0}+\sqrt{1-x_{i}^{2}}\ket{1}$ and the feedback rule
$R_{t}=R_{t-1}+\frac{p_{(t-1)2}-R_{t-1}}{m},$ (4)
where $p_{ti}$ is the probability of detecting one photon in the $i$-th mode ($i=0,1,2$), at time $t$, while $m$ is the fixed memory extent of the memristor (see section IB of the Supplemental Material for further details).
For both tasks, the collected experimental statistics are fed into the classical readout ANN that is trained to perform the given task.
As a benchmark, we performed the task also without any feedback loop and with no reservoir. The latter is a fully classical procedure, only injecting the x values into the final ANN, to highlight the role of our physical platform.
III Experimental implementation
The main part of the experimental setup, fully represented in Fig. 2, is the physical quantum reservoir for information processing. It consists of an integrated waveguide circuit with 3 input/output modes and is designed according to a universal architecture [44]. The waveguides are fabricated by direct femtosecond laser writing [35, 45, 46] in an alumino-borosilicate glass substrate and feature single-mode operation at 1550 nm. Microstructured thermo-optic phase shifters are integrated on-chip, to provide efficient phase tunability [47]. Input and output ports are pigtailed with single-mode fibers.
The encoding of the classical input is set via an initial unitary transformation $\textbf{U}_{enc}$, through an on-chip MZI, as shown in Fig. 1b. For example, to encode the state $\sqrt{x}|0\rangle+\sqrt{1-x}|1\rangle$, apply a phase $\phi=2\arccos(x)$ when injecting a photon into the lower mode of the MZI. However, since the PIC consists of three MZIs in total, we implement the product $\textbf{U}_{1}\cdot\textbf{U}_{enc}$ through the first MZI, as detailed in Fig. 2b.
The input single photon state is generated by a collinear Type-II spontaneous parametric down conversion source, pumped at 775 nm, which emits pairs of degenerate photons with a wavelength of 1550 nm. The employed non-linear crystal is a periodically poled titanyl phosphate, placed within a Sagnac interferometer. Since for this application we do not need entangled states, the crystal is pumped only in one spatial direction and generates the separable state $|01\rangle$. One photon is ancillary and heralds the generation of the second photon that is injected into the photonic processor. The three output modes of the photonic processor are connected to superconducting nanowire single photon detectors.
At this point, the frequencies of the clicks registered by the three detectors at the output of the chip, in coincidence with the one detecting the heralding photon, give the probabilities $p_{0}$, $p_{1}$ and $p_{2}$ as $p_{i}=\frac{N_{i,a}}{\sum_{j=1}^{3}N_{j,a}}$.
IV Results
To quantify the quality of the performance of predictive tasks, we use the so-called mean squared error [48], defined as follows:
$MSE(\bm{y},\tilde{\bm{y}})\coloneq\frac{1}{N}\sum_{t=1}^{N}\left(y_{t}-\tilde{y}_{t}\right)^{2}$ (5)
with $y_{t}$ target values, $\tilde{y}_{t}$ predicted values and $N$ data points.
For the function prediction task three different cases are compared: (i) with quantum memristor, (ii) without memory (with no feedback loop) and (iii) with no reservoir. The prediction of $x^{n}$ with $n=4$ is shown in Fig. 3 and the inset plot shows the loss function of the readout unit for different $n$. Note that even without the feedback loop some non-linearity can be generated, due to the encoding from $x$ to $\rho_{x}$. However, the feedback loop enhances the impact of the non-linear measurement process and ultimately the performance of the algorithm. On the contrary, in the case of no reservoir, the output is simply a linear function, as the final ANN may only act linearly on the input.
MODEL EQUATION MSE (×$\times 10^{-4}$)
$L$ $y_{t}\approx\alpha u_{t}+\beta$ $2.76(29)$
$Q$ $y_{t}\approx\alpha u_{t}^{2}+\beta u_{t}+\gamma$ $2.05(22)$
$L+M$ $y_{t}\approx\alpha u_{t}+\beta u_{t-1}+\gamma$ $1.83(19)$
$Q+M$ $y_{t}\approx\alpha u_{t}^{2}+\beta u_{t-1}+\gamma$ $1.22(14)$
$L^{M}$ $y_{t}\approx\alpha u_{t}+\beta\;\text{Exp}_{4}(u_{t})+\gamma$ $1.14(11)$
$Q^{M}$ $y_{t}\approx\alpha u_{t}^{2}+\beta\;\text{Exp}_{4}(u_{t})+\gamma$ $0.78(8)$
QMEM - $\bm{0.33(8)}$
Table 1: Mean Squared Errors for classical machine learning models. This table shows the average MSE achieved by several classical models, in comparison to the one presented in this work, indicated by QMEM. The numbers in the bracket indicate one standard deviation of 100 runs. In detail, $L$ and $Q$ stand, respectively, for models that can base their predictions on linear and quadratic manipulation of the input. $L+M$ ($Q+M$) indicates the same models, but with access to the previous element of the output series. In $L^{M}$ ($Q^{M}$), instead, the access to previous elements is extended to the exponential moving average of all the past outputs, with a memory extent of 4. The reported values are obtained through numerical simulations.
In Fig. 4, we show the experimental prediction of the network for the previously described NARMA task. For the case with no feedback loop, each output depends only on the last input, since only the feedback loop allows storing the information of previous data points. It is noteworthy that the case with no feedback loop still has decent performance, due to the low non-linearity of the task (see section IC of the Supplemental Material for further details). To have a more complete overview of the performance of our algorithm, we compared it with similar classical algorithms. In particular, we tested the aforementioned tasks with models that could base their predictions on a linear or quadratic manipulation of the inputs in three cases: with no memory, with access to the previous outcome and to the exponential moving average of all past inputs, with memory extent equal to 4. For all these cases, the average MSE obtained in our case is lower, highlighting the enhancement brought by the quantum memristor (the numerical values are reported in Table I).
V Discussion
In this work, we have implemented the first instance of QRC exploiting single photon states based on a quantum memristor device. These machine learning models work by injecting input data into a fixed random reservoir, which is then followed by a one-layered linear ANN (the only part of the model which is trained) [34]. The required non-linearity is provided by the evolution of the quantum system through the photonic quantum memristor.
The last part of the model, featuring an ANN, is implemented through standard computing. We tested our model on two function prediction tasks: several power functions and the NARMA task. Due to its regularity, the first case is only to test the non-linearity that is achievable through our apparatus. The latter instead tests also the memory of our model, given that its outcomes depend on previous ones.
Ours is the first example of neuromorphic architectures exploiting single-photon states (encoded in the path degree of freedom) and implementing a non-linear behaviour through an adaptive protocol, performed by the quantum memristor. Moreover, we carry out a systematic analysis to distinguish the source of such a non-linearity, i.e. the encoding or the feedback loop. For this reason, for all of the tasks, we compare the case featuring the quantum memristor (QRC), to the case without feedback loop. For all of the cases, the quantum memristor brings an enhanced accuracy with respect to classical models. This scheme, although applied to relatively simple tasks, can be seen as a building block for more complex networks, as the output of the memristor, tracing out the update mode, preserves coherence and can be used as an input for successive layers (see section II of the Supplemental Material for further details).
This study paves the way for further investigations related to the non-linearities achievable on photonic platforms. This is particularly relevant, considering that hybrid optical-electronic ANNs have been proven to require a lower energy consumption than standard ones [25, 26].
In this framework, the memristor could be employed to constitute the activation function layer of quantum neural networks. Another interesting feature is that a quantum reservoir could be used to directly process quantum systems, tackling tasks with no classical equivalent. For instance, to investigate foundational aspects or quantum state properties [49, 50, 51]. Further studies can involve the implementation of a quantum feedback, where previous states interfere with new ones, through a loop architecture [52]. Analogously, the implementation of a fast feedback would allow to make coherent adjustments to the quantum state. This would allow for a larger memory storage and open up to the possibility of tackling a wider range of tasks.
Acknowledgments
The fabrication of the photonic integrated circuit was partially carried out at PoliFAB (https://www.polifab.polimi.it/), the micro- and nano-fabrication facility of Politecnico di Milano. F.C. and R.O. would like to thank the PoliFAB staff for technical advice and support. This research was funded in whole or in part by the Austrian Science Fund (FWF)[10.55776/ESP205] (PREQUrSOR), [10.55776/F71] (BeyondC), [10.55776/FG5] (Research Group 5) and [10.55776/I6002] (PhoMemtor).
Competing interests
The authors declare no competing interest.