### v2

# Experimental neuromorphic computing based on quantum memristor

Abstract
Machine learning has recently developed novel approaches, mimicking the synapses of the human brain to achieve similarly efficient learning strategies. Such an approach retains the universality of standard methods, while attempting to circumvent their excessive requirements, which hinder their scalability. In this landscape, quantum (or quantum inspired) algorithms may bring enhancement. However, high-performing neural networks invariably display nonlinear behaviours, which poses a challenge to quantum platforms, given the intrinsically linear evolution of closed systems. We propose a strategy to enhance the nonlinearity achievable in this context, without resorting to entangling gates and report the first neuromorphic architecture based on a photonic quantum memristor. In detail, we show how the memristive feedback loop enhances the nonlinearity and hence the performance of the tested algorithms. We benchmark our model on four tasks, a nonlinear function and three time series prediction. In these cases, we highlight the essential role of the quantum memristive element and demonstrate the possibility of using it as a building block in more sophisticated networks.
I Introduction
In the last decades, machine learning has fundamentally changed the approach to solving a wide variety of problems, from everyday life to scientific scopes [1, 2, 3, 4, 5, 3, 6, 7]. In this context, artificial neural networks (ANN) have proven to be a pivotal tool, especially for tackling tasks where information needs to be extracted from high-dimensional data [8]. Nevertheless, a growing number of parameters must be trained as the tasks get more complex, with correspondingly larger training data, resulting in the so-called memory bottleneck [9]. This sparked the development of more dedicated hardware, i.e. neuromorphic architectures, which mimic the human brain [10]. The main advantage of these models is their minimal learning requirements, which circumvents the aforementioned difficulty. Acting as the poster-child of neuromorphic components is the memristor, a passive circuital component that was postulated in 1971 [11] and demonstrated in 2008 [12]. Its main characteristic is that it retains memory of its past states, in the form of a hysteresis that is similar to neural synapses [13, 14, 10, 15, 16].
Figure 1: Neuromorphic computing based on quantum memristor. The basic principle of a quantum memristor is a tunable Mach-Zehnder interferometer, where its internal phase ($\theta$) is varied, according to the output registered at one of its output modes, through a feedback loop. The investigated hybrid quantum/classical models are implemented through three parts: input encoding (quantum), nonlinear reservoir (quantum) and linear regression, i.e. $y_{out}=\Sigma y_{i}w_{i}$ (classical. In our work, the encoding of classical variables in a one qubit state exploits the amplitude of the component parallel to $|0\rangle$, i.e. $x\rightarrow\sqrt{x}|0\rangle+\sqrt{1-x}|1\rangle$ and $x|0\rangle+\sqrt{1-x^{2}}|1\rangle$. The reservoir consists in a unitary operation, composed by the product of different rotations. These rotations allow us to achieve a nonlinear transformation of the input, due to the encoding that is nonlinear in the density matrix of the quantum state. The loop enhances the achievable nonlinearity, while implementing a short-term memory. The nonlinear reservoir is then followed by a linear regression model, which is the only part of the model that is trained, producing a weighted sum of the reservoir outputs. The four addressed tasks are the prediction of a smooth nonlinear function (FP) and three random time series: NARMA (N), Mackey-Glass (MG), and Santa Fe (SF).
Figure 2: Experimental setup. (a) The experimental apparatus includes a source of single photon pairs pumped with a laser at 775 nm and generates degenerate photon pairs at 1550 nm. One photon is routed to the 3 input/output mode photonic integrated circuit (PIC) and the other photon is ancillary. Let us note that this circuit is implemented following the universal Reck architecture [17]. At the outputs of the PIC, photon counts are registered by superconducting nanowire single photon detectors. After the full calibration of the chip, phases are set and modulated by applying a suitable voltage to the resistive elements, that act as tunable thermal phase shifters. The unitary operation is updated, conditioned on the statistics of previous runs. The output distribution is then fed into a linear regression model, which is the only part of the machine learning model that is trained. (b) The optical circuit consists in 3 Mach-Zehnder interferometers and 5 tunable phases. The $R_{i}$ indicate the resistive elements, where $R_{3}$ is updated conditioned on the output of the third mode, implementing a memristive behaviour. Let us note that the state obtained tracing out the third output still preserves coherence (see Supplementary Material note I). The input is encoded in one qubit, through a unitary transformation $\mathbf{U}_{\text{enc}}$. The unitary transformations $\mathbf{U}_{1}$ and $\mathbf{U}_{2}$ are taken as hyperparameters of our model for the monomial functions. Note that we implement the product $\mathbf{U}_{1}\cdot\mathbf{U}_{\text{enc}}$ with the first MZI. Then, $\mathbf{U}_{\text{mem}}$ amounts to the action of the memristor and is varied depending on previous outcomes. Also the coefficient of the linear feedback rule, according to which this happens, are hyperparameters, see Eq. (2). Let us note that for the time series prediction tasks, $U_{1}=U_{3}=\mathbb{I}$, and the only hyperparameter is constituted by the memory decay $m_{d}$ (see Eq. (4)).
Quantum computing introduced a related but somewhat orthogonal paradigm shift in computer science. The interest in this field stems, first, from the possibility of tackling non-classical processing tasks directly related to the investigation of quantum effects and, second, from the promise of outperforming standard algorithms for particular problems [18, 19, 20, 21]. However, for the latter, the only proven advantages were shown for tasks that are beyond the reach of even near-term state-of-the-art quantum computers or which have no known applications [22]. Hence, a research line that has attracted a lot of interest arises from the combination of machine learning and quantum computation, that is, quantum machine learning (QML) [23, 24]. This is fueled by the hope that a computational advantage can be found on practically relevant tasks even when using relatively small-scale quantum information processing. In this context, photonic apparatuses constitute a very promising platform, as they exhibit a vastly lower energy consumption compared to standard electrical ones [25, 26], indicating the possibility of far more efficient and fast computational platforms. However, when considering quantum processes, the natural way of implementing nonlinear evolutions (crucial for machine learning algorithms) consists in the (controlled) interaction among multiple systems, i.e. entangling gates. This kind of evolution is still governed by linear transformations, but it can be seen as nonlinear from a computational point of view. These can be deterministically achieved on several platforms, e.g. superconducting qubits or trapped ions, but the scalability is hindered by the rising of decoherence. In contrast, on photonic platforms, decoherence is less of an issue, but photonic nonlinear interactions are weak and hardly controllable. This implies, for example, that very high optical powers need to be used, which, besides the engineering challenges, can also negate the potential energy advantage coming from using optics instead of electronics. A possible way out is by resorting to optoelectronics [26], although it requires energy to convert light into electronic signals and back [25], and it slows down the whole process. This is the reason why, until now, this problem has been circumvented, either through a suitable encoding of the data, as in [27, 28], or by exploiting the nonlinearity coming from the measurement, as in [29, 30]. These approaches are all proven to be effective, but they present drawbacks. Indeed, the first cannot be used with quantum data and the second cannot implement any memory nor be used as a subroutine.
In this work, we propose a possible approach to solve this apparent conundrum and we implement the first instances of a neuromorphic computing protocol, i.e. reservoir computing (RC), using single photons, manipulated through an integrated circuit [31]. The model (see Fig. 1) may be seen as a building block for more complex networks and is composed of a nonlinear fixed reservoir (equipped with classical memory), followed by the readout unit, implementing a linear regression model, $y_{out}=\Sigma y_{i}w_{i}$, which is the only part that requires a training phase. Our aim is to implement a quantum/classical hybrid algorithm, where the nonlinearity is generated by a quantum system, while the linear regression model is classical. To achieve this goal we leverage a particular choice for the input encoding and the exploitation of a novel device: the photonic quantum memristor [32]. This element has the two-fold effect of enhancing the nonlinearity of the model, while implementing a classical memory. This device is based on a Mach-Zehnder interferometer equipped with a feedback loop, and its behaviour is proven to be memristive. Moreover, the impact of the feedback on the coherence of the output state was studied, both in a configuration featuring a single device and in a network of two independent quantum memristors, injected of correlated photon pairs [33]. However, its potential use for machine learning tasks was only theorized and its role was not deeply investigated, as the task proposed in the original manuscript, i.e. the recognition of MNIST hand-written digits [34], can be tackled effectively through linear regression and other simple estimators. Before scaling up to larger systems, we investigate the potentialities of one device, to understand what the minimal requirements are to tackle real-world machine learning tasks. To give a better estimation, we also carry out a comparison between our model and simple non-task-specific classical models, to quantify the resources they require to achieve a similar, or slightly better, performance. Hence, here, we go beyond the previous study and apply this hardware to four machine learning tasks where nonlinearity is a crucial factor and cannot be omitted without compromising performance, highlighting the pivotal role of the quantum reservoir. In all of the considered tasks, we show that the dynamics of the photonic quantum memristor enhances the performance, with respect to the case where no feedback loop is implemented. Hence, the present work shows how a (photonic) quantum system can tackle tasks that were addressed with similar classical architectures [14], but with substantially less resources, paving the way to resource efficient machine learning models and to the possibility of exploring neuromorphic architectures applied to quantum tasks.
II Quantum memristor-based neuromorphic models
The photonic quantum memristor can be modeled as a tunable Mach-Zehnder interferometer, as depicted in Fig. 2. Then, its internal phase is updated according to a feedback rule that depends on the measurement outcome at one output mode, affecting the temporal dynamics of the device. One of its main features consists in the fact that, when tracing out the mode used for the feedback loop, the remaining reduced state may still be partially coherent relative to the joint state, as detailed in the Supplementary Material note I. Consequently, further quantum information processing is possible on the output.
We want to exploit this feedback mechanism to implement nonlinear operations on the input state and utilize a short-term memory, to achieve a proof-of-principle quantum reservoir computing (QRC). In detail, we encode our classical inputs in quantum states that are then injected into a quantum reservoir. This usually consists of randomly connected nodes, whose internal time-based state make up its intrinsic memory and the historical nonlinearity that is necessary for the learning process (see the Supplementary Material note II for further details). In this framework, feedback-driven and (weak-) measurement-based architectures have been recently theoretically investigated [35, 36, 37]. Our implementation features one physical node, realized through a Mach-Zehnder interferometer (MZI), which in concert with a feedback loop acts like a quantum memristor. Then, the randomness is given by unitary operations before and after the memristor itself. Hence, the quantum part of our model consists in the encoding of the classical data and reservoir. In particular, we exploit one-qubit states, encoded in the path degree of freedom of a single photon state, with a dual-rail scheme (see Fig.  2b). The reservoir is then composed of a sequence of three unitaries over three spatial modes, i.e. $\textbf{U}_{1}$, $\textbf{U}_{mem}$ and $\textbf{U}_{2}$. At the end, to build a proper reservoir computing scheme, a readout unit, i.e. a classical linear regression model, follows.
Figure 3: Nonlinear function prediction results. The blue dots show the experimental results for the prediction of the power function $f_{t}(x)=x^{n}$ for $n=4$ (indicated by the solid black curve). The orange ones show the results for the same task, but without feedback loop. Inputs lower than 0.9 are used for training the model, while the other ones are used for test. The case with no feedback loop is on average 34% less accurate than the one exploiting the memristor. Let us note that, to have a fair comparison, hyperparameter were optimized separately for the two cases. The dashed black line represents the case without the physical reservoir, i.e. where the inputs $x$ are directly injected to the readout linear regression. The loss function is defined as the mean squared error MSE between the target function $f_{t}(x)=y_{t}$ and the test data $y^{k}_{test}$, i.e. $y_{t}-y_{test}$. The inset shows the MSE for power function with $n=3,4,5,6$ in the QRC case (blue) and with no feedback loop case (orange).
We test our model for four tasks, that rely on the two key features of our photonic platform, i.e. nonlinearity and short-term memory. These are a nonlinear function prediction/reproduction, where we exploit mostly the nonlinearity of the model, and three different time series predictions nonlinear Auto Regressive Moving Average, i.e. NARMA, [38, 39] Mackey-Glass dynamic prediction [40, 41] and SantaFe laser intensity prediction [42, 43], where short-term memory plays a pivotal role.
For the first task, the functions we aim at reproducing and then extrapolating are monomials:
$f(x)=x^{n},\ n\in\mathbb{Z}$ (1)
This choice is motivated by the interest in testing the maximal achievable nonlinear behaviour of our model and the role of the feedback loop towards its enhancement.
We encode the inputs as follows: $x\rightarrow\sqrt{x}|0\rangle+\sqrt{1-x}|1\rangle$. Note that this encoding is linear when performing a projective measurement in the eigenbasis of the Pauli operator $\sigma_{Z}$ [44], as $Tr(\sigma_{Z}\rho_{x})=\sum_{i=0,1}(-1)^{i}p_{i}=2x-1$, where $p_{i}$ is the probability of getting the outcome $i$. In contrast, measurements in the $\sigma_{X}$ basis give an expectation value that is nonlinear in $x$, $Tr(\sigma_{X}\rho_{x})=\sqrt{x}\sqrt{1-x}$. This nonlinearity can be enhanced by a proper choice of the rotations $\textbf{U}_{1}$ and $\textbf{U}_{2}$. Then, we pick a linear feedback rule which modifies $\textbf{U}_{mem}$. The feedback loop additionally enhances the nonlinearity, by inducing a dependence of the measurement operator on the previous output state. This results in a nonlinear form of the output, i.e. $(x_{t-1}x_{t})$, already after one step (for the full analytical derivation, see Supplementary Material note II), and also in a memory of the previous states. From an operational point of view, the internal variable of the memristor $R_{t}$, which can be seen as the reflectivity of a beam-splitter, is updated via:
$R_{t}=\frac{1}{m}\sum_{T=t-m}^{t}R_{T}=\frac{1}{m}\sum_{T=t-m}^{t}(a\cdot p_{T,2}+b)$ (2)
where $p_{T,2}$ is the probability of detecting a photon in the update mode and the internal phase of the MZI that implements $\textbf{U}_{mem}$ is given by $2\times arccos(\sqrt{R})$. The rotations $\textbf{U}_{1}$ and $\textbf{U}_{2}$, along with $a$ and $b$, are optimizable hyperparameters (see Supplementary Material note IIA for further details). In this case, $m$ can be seen as the memory extent, which amounts to the number of previous steps which are used when performing the average reported in Eq. (2), to retrieve the reflectivity at each time step $t$.
The goal of the second task is to predict the output of a nonlinear dynamical system, introduced in [39, 38], called NARMA. This was already widely used as a benchmark for classical [45, 46] and quantum [41, 47] RC models. The recurrence relation of the system is described by the following equation:
$y_{t+1}=0.4y_{t}+0.4y_{t}y_{t-1}+0.6x_{t}^{3}+0.1,$ (3)
where $x_{t}$ and $y_{t}$ is the input and the output, respectively, at time $t$. From Eq. (3), it is visible that the output $y_{t+1}$ depends on past two outputs $y_{t}$ and $y_{t-1}$.
The time series $\bm{x}=\{x_{t}\}_{t=1}^{N}$ is generated by sampling each $x_{t}$ independently and uniformly in the interval $U(0,\frac{1}{2})$, and then the outputs $\bm{y}=\{y_{t}\}_{t=1}^{N}$ come sequentially from Eq. (3). For this task, we are using the amplitude encoding $x_{t}\mapsto x_{t}\ket{0}+\sqrt{1-x_{t}^{2}}\ket{1}$ and the feedback rule
$R_{t}=R_{t-1}+\frac{p_{t-1,2}-R_{t-1}}{m_{d}},$ (4)
where $p_{ti}$ at time $t$ is the probability of detecting one photon in the $i$-th mode ($i=0,1,2$), while $m_{d}$ is the fixed memory decay of the memristor. The memory decay, analogously to the memory extent mentioned before, is also a hyperparameter, but it is related to how fast the memory of past outcomes will decay. Indeed, this feedback rule does not involve any average and, as detailed in the Supplementary Material note IIB, it modulates the dependence of $R_{t}$ on each previous output.
The other two tasks are one-step-ahead predictions of a real data series. This implies that, given $y_{t}$, the model needs to predict $y_{t+1}$. The first of them is the solution to the Mackey-Glass differential equation [40, 41] and the last is the so-called Santa Fe time series, which is a dataset recorded from a far-infrared laser in a chaotic state [42, 43]. For both tasks, we use amplitude encoding $x_{i}\mapsto\sqrt{1-x_{i}}\ket{0}+\sqrt{x_{i}}\ket{1}$ and the feedback rule in Eq. (4) with $m_{d}=2$ for Mackey-Glass task and $m_{d}=6$ for Santa Fe task. Both data sets have been scaled down to fall within the range of 0 to 1. For more details see the Supplementary Material note IIB.
For all the tasks, after the quantum processing, a classical linear regression is used to find the best prediction. We note that during this process no nonlinearity or memory is added. As a benchmark, we performed the monomial prediction tasks without any feedback loop and with no reservoir. The latter is a fully classical procedure, only injecting the x values into the final linear regression, to highlight the role of our physical platform. For the time series predictions, instead, we performed a comparison with slightly more complex fully classical models, even featuring some nonlinearity and memory of past states.
Figure 4: NARMA, Mackey-Glass, and Santa Fe task results. We show the performance of our model on the following prediction tasks: NARMA (A), Mackey-Glass (B), and Santa Fe (C). The true values of the time series are plotted in black. Predictions with and without the memristor are shown, respectively, in blue and orange. In each subplot, the MSE score is reported in the top right corner. Each of the adopted datasets contains 1000 points. Out of those, the first 20 correspond to the so-called washout, i.e. the time required to the model to initialize. Then, the training process is carried out on the following 480 points, while the validation is performed on the rest of the dataset. For the sake of clarity, the plot shows only 200 points of the test.
III Experimental implementation
The main part of the experimental setup, fully represented in Fig. 2, is the physical quantum reservoir for information processing. It consists of an integrated waveguide circuit with 3 input/output modes and is designed according to a universal architecture [48]. The waveguides are fabricated by direct femtosecond laser writing [32, 49, 50] in an alumino-borosilicate glass substrate and feature single-mode operation at 1550 nm. Microstructured thermo-optic phase shifters are integrated on-chip, to provide efficient phase tunability [51]. Let us highlight that, due to the limited size, our circuit does not suffer from thermal crosstalk (for more detailed information about the calibration procedure, see the Supplementary Material note III). Furthermore, due to the low birefringence of the substrate, this circuit is polarization independent. Input and output ports are pigtailed with single-mode fibers.
The encoding of the classical input is set via an initial unitary transformation $\textbf{U}_{enc}$, through an on-chip MZI, as shown in Fig. 1b. For example, to encode the state $\sqrt{x}|0\rangle+\sqrt{1-x}|1\rangle$, apply a phase $\phi=2\arccos(x)$ when injecting a photon into the lower mode of the MZI. However, since the optical circuit consists of three MZIs in total, we implement the product $\textbf{U}_{1}\cdot\textbf{U}_{enc}$ through the first MZI, as detailed in Fig. 2b.
The input single photon state is generated by a collinear Type-II spontaneous parametric down conversion source, pumped at 775 nm, which emits pairs of degenerate photons with a wavelength of 1550 nm. The employed nonlinear crystal is a periodically poled titanyl phosphate, placed within a Sagnac interferometer. Since for this application we do not need entangled states, the crystal is pumped only in one spatial direction and generates the separable state $|01\rangle$. One photon is ancillary and heralds the generation of the second photon that is injected into the photonic processor. This implies that we post-select our statistics, considering only the cases where both the heralding and signal photons are detected. The three output modes of the photonic processor are connected to superconducting nanowire single photon detectors.
At this point, the frequencies of the clicks registered by the three detectors at the output of the chip, in coincidence with the one detecting the heralding photon, give the probabilities $p_{0}$, $p_{1}$ and $p_{2}$ as $p_{i}=\frac{N_{i,a}}{\sum_{j=1}^{3}N_{j,a}}$.
IV Results
To quantify the quality of the performance of predictive tasks, we use the so-called mean squared error [52], defined as follows:
$MSE(\bm{y},\tilde{\bm{y}})\coloneq\frac{1}{N}\sum_{t=1}^{N}\left(y_{t}-\tilde{y}_{t}\right)^{2}$ (5)
with $y_{t}$ target values, $\tilde{y}_{t}$ predicted values and $N$ data points.
For the monomial prediction tasks three different cases are compared: (i) with quantum memristor, (ii) without memory (with no feedback loop) and (iii) with no reservoir. The prediction of $x^{n}$ with $n=4$ is shown in Fig. 3 and the inset plot shows the loss function of the readout unit for different $n$. Note that even without the feedback loop some nonlinearity can be generated, due to the encoding from $x$ to $\rho_{x}$. However, the feedback loop enhances the impact of the nonlinear measurement process and ultimately the performance of the algorithm. Let us highlight that, when comparing the cases with and without feedback, we are picking the best hyperparameters (namely $U_{1}$ and $U_{3}$ and the coefficients of the feedback function), separately for the two cases, to have a fair comparison. In contrast, in the case of no reservoir, the output is simply a linear function, since the final linear regression only gives a weighted sum of the inputs.
MODEL EQUATION MSE (×$\times 10^{-4}$)
$L$ $y_{t}\approx\text{Poly}_{1}(x_{t})$ $2.76(29)$
$C$ $y_{t}\approx\text{Poly}_{3}(x_{t})$ $2.05(22)$
$L+M$ $y_{t}\approx\text{Poly}_{1}(x_{t},x_{t-1})$ $1.83(19)$
$C+M$ $y_{t}\approx\text{Poly}_{3}(x_{t},x_{t-1})$ $0.92(11)$
QMEM - $\bm{0.33(8)}$
Table 1: Mean Squared Errors for classical machine learning models for NARMA task. This table shows the average MSE achieved by several classical models, which have access to a similar amount of resources, in terms of memory and free variables, in comparison to the one presented in this work, indicated by QMEM. The numbers in the bracket indicate one standard deviation of 100 runs. $\text{Poly}_{d}(a_{1},\ldots,a_{n})$ is the polynomial of variables $\{a_{1},\ldots,a_{n}\}$ of degree at most $d$. In detail, $L$ and $C$ stand, respectively, for models that can base their predictions on linear and cubic manipulation of the input. The first features 2 and the second 4 free variables. $L+M$ ($C+M$) indicates the same models, but with access to the previous element of the output series. In this case, the first features 3 and the second 9 free variables. For an accurate description of such models, see Supplementary Material note II. The reported values are obtained through numerical simulations. We note that polynomials of higher order improve neither of the results. Our model features only 1 free variable, given by the memory decay $m_{d}$ (see Eq. (4)).
In Fig. 4, we show the experimental prediction of the network for the previously described time series prediction tasks. For the case with no feedback loop, each output depends only on the last input, since there is no storing the information of previous data points. It is worth noting that cases without feedback loops still achieve fairly good results (although worse than the case where the quantum memristor is employed). For the NARMA task, this is due to its low nonlinearity, while for the Mackey-Glass and Santa Fe tasks, it results from the high density of probing points, which makes the prediction almost linear. To have a more complete overview of the performance of our algorithm, we present the results achieved on the same tasks by simple classical algorithms. First, we consider a model without memory, where the prediction is a polynomial function of the input data. In particular, we consider two models, featuring 2 and 4 free parameters. Next, we examine two models with the simplest possible memory, which records only the previous input. Hence, the final prediction will be given by a polynomial function of the current and previous input. In this case, the two models have, respectively, 3 and 9 free parameters. Our model featuring the quantum memristor has a slightly better performance than those models (see Table 1), even when the classical one has access to past inputs. For the comparison of the other tasks and the mathematical description of these models, see Supplementary Material note IIC.
V Discussion
In this work, we have implemented the first instance of QRC exploiting single photon states based on a quantum memristor device. These machine learning models work by injecting input data into a fixed random quantum reservoir, implemented through a photonic quantum memristor, which is then followed by a classical linear regression model [31]. We tested our algorithm on four function prediction tasks: several monomial functions, and then NARMA, Santa Fe and Mackey-Glass time series. Due to its regularity, the first case is only to test the nonlinearity that is achievable through our apparatus. The others, instead, test also the memory of our model, given that its outcomes depend on previous ones.
Ours is the first example of neuromorphic architectures exploiting single-photon states (encoded in the path degree of freedom) and implementing a nonlinear behaviour through an adaptive protocol, performed by the quantum memristor. Moreover, we carry out a systematic analysis to distinguish the source of such a nonlinearity, i.e. the encoding or the feedback loop. For this reason, for all of the tasks, we compare the case featuring the quantum memristor (QRC), to the case without feedback loop, consistently showing that the quantum memristor brings an enhanced accuracy with respect to classical models. This scheme, although applied to relatively simple tasks, can be seen as a building block for more complex networks, as the output of the memristor, when tracing out the update mode, preserves coherence and can be used as an input for successive layers (see the Supplementary Material note I for further details).
A potential way of enhancing the expressivity of the proposed model is to use multiple memristors in a consecutive way, as shown for instance in [53], and updating both the internal variables of those memristors, as well as the input state, so that new inputs would interfere with states encoding previous outcomes. Analogously, the implementation of a fast feedback would allow to make coherent adjustments to the quantum state. This would result in a larger memory storage and open up to the possibility of tackling a wider range of tasks. Moreover, feeding parts of the outputs as new inputs, for instance through a loop architecture [54], could unlock the possibility of tackling forecasting tasks [55]. Another interesting feature is that a quantum reservoir could be used to directly process quantum systems, tackling tasks with no classical equivalent. For instance, to investigate foundational aspects or quantum state properties [29, 56, 57].
Besides exploring the possibilities unlocked by the quantum features of the memristor, this study paves the way for further investigations related to the nonlinearities achievable on photonic platforms, for optical computing. This is particularly relevant, considering that hybrid optical-electronic ANNs have been proven to require a lower energy consumption than standard ones [25, 26]. In this framework, the memristor could be employed to constitute the activation function layer of quantum neural networks and, after further investigations on the relation between the memristor’s dynamics with neurons one, it could unlock the possibility of implementing spiking optical neural networks [58, 59, 60].
Acknowledgments
The fabrication of the photonic integrated circuit was partially carried out at PoliFAB (https://www.polifab.polimi.it/), the micro- and nano-fabrication facility of Politecnico di Milano. F.C. and R.O. would like to thank the PoliFAB staff for technical advice and support. This research was funded in whole or in part by the Austrian Science Fund (FWF)[10.55776/ESP205] (PREQUrSOR), [10.55776/F71] (BeyondC), [10.55776/FG5] (Research Group 5) and [10.55776/I6002] (PhoMemtor). R.O. acknowledges financial support by ICSC – National Research Center in High Performance Computing, Big Data and Quantum Computing, funded by European Union– NextGenerationEU.
Author Contribution
Mir.S. and I.A. designed and conducted the experiment, Mich.S. developed the theory and algorithm for the time series predictions and analyzed the corresponding data, I.A. developed the algorithm for power predictions and I.A. and Mir.S. analyzed the corresponding data. R.A., A.C. and F.C. conducted the design, fabrication and calibration of the integrated photonic processor. Mir.S., I.A. and Mich.S. wrote the first draft of the manuscript. I.A., Mag.S., R.O., B.D. and P.W. supervised the whole project. All authors discussed the results and reviewed the manuscript.
Competing interests
F.C. and R.O. are co-founders of the company Ephos. The authors declare that they have no other competing interests.